# Uncertainty Quantification for LLMs through Minimum Bayes Risk: Bridging Confidence and Consistency

### v4

@@ Abstract
Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompass a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling. Several recent methods have combined these two approaches to boost UQ performance. However, they sometimes fail to outperform much simpler baseline methods. Our work discusses the fundamental approach to constructing uncertainty measures that directly links uncertainty with the minimum Bayes risks achieved by LLM decoding. Building on these findings, we propose a novel approach to integrating model confidence with output consistency, resulting in a family of efficient and robust UQ methods. We evaluate our approach across various tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advances in information retrieval Zhu et al. (2023), question answering Kwiatkowski et al. (2019), machine translation Kocmi and Federmann (2023), and a broad range of other NLP applications. As these models become an integral part of our everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which estimates the reliability of model outputs and enables abstention from generations that are likely to be incorrect.
@@
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (Malinin and Gales, 2021; Fomicheva et al., 2020). A key limitation of these methods is that they cannot account for semantic variability across multiple possible responses for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationship between the sampled outputs (Fomicheva et al., 2020; Lin et al., 2024), capturing uncertainty as objective variability of meaning among the sampled outputs.
Information-theoretic and consistency-based methods have complementary strengths. For this reason, recent state-of-the-art methods aimed to unify these approaches (Kuhn et al., 2023; Duan et al., 2024). We follow this direction and propose a way of quantifying risk as a combination of these basic measures. This results in a family of efficient and robust UQ techniques. Our approach combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Recently, it has been argued Wang and Holmes (2025); Daheim et al. (2025) that UQ in NLP tasks can be viewed through the lens of the Minimum Bayes Risk (MBR) framework. Following this, we formulate our methods as particular forms of risk functions under the MBR approach.
@@
We propose a new way of quantifying risk that combines information-theoretic and consistency-based measures, and derive a family of Confidence and Consistency-based Approaches (CoCoA) to uncertainty quantification in LLMs.
We show that these measures can be approximated by a learned function trained on unlabeled held-out set with a negligible loss of performance, eliminating the need for costly repeated sampling from the LLM.
We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and machine translation. Our experiments demonstrate sizable improvements in the reliability and robustness of UQ compared to state-of-the-art methods.
@@ 2.1 Language Model Decoding
LLMs define a probabilistic output distribution for the given input sequence . The standard ways to obtain a particular output given include various variants of sampling and greedy decoding, where
An alternative approach to decoding is Minimum Bayes Risk (MBR) decoding Kumar and Byrne (2004):
@@
with being a pairwise loss function. The standard choice in the MBR literature is to take , where the so-called utility function represents some notion of similarity between generated sequences and . Below we discuss how various decoding strategies and corresponding risk functions lead to well-grounded definitions of uncertainty.
@@ 2.2 From Risk to Uncertainty
Bayesian uncertainty measures are well known to be strongly connected with minimum risks of various kinds Xu and Raginsky (2022); Kotelevskii et al. (2025). Recently, this connection was revisited in the context of natural language generation Wang and Holmes (2025); Daheim et al. (2025), leading to the development of new MBR-based UQ methods. More precisely, one can consider the smallest achievable risk within or, equivalently, to maximum expected utility:
@@
Single Sequence Information-Based Methods. Information-based methods estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. Within the framework of MBR decoding, consider . In this case, Bayes risk corresponds to the expected zero-one loss of decoding:
This leads to one of the simplest information-based uncertainty measures, Maximum Sequence Probability (MSP):
Although widely used as an uncertainty measure across various applications, including natural language generation, its connection to statistical risk has only recently been explored in the UQ literature Kotelevskii et al. (2022, 2025); Aichberger et al. (2024).
Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix D.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
@@
LLMs provide us only the probability of a specific answer, even though the same meaning can often be conveyed in multiple ways. To obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
@@
Reliability of such methods depends heavily on the underlying LLM being well-calibrated, which is a quality that is hard to define and harder to ensure.
@@
Semantic Consistency-Based Methods. The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs , where . Consistency-based UQ methods rely on the diversity of answers sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require a probability distribution estimated by LLM and can be applied in the black-box setting, when only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, consistency-based method start from defining some similarity function between arbitrary generations and . The value indicates the complete equivalence between and , and indicates that there is no similarity. Similarity could be computed in various ways. For instance, the Lexical Similarity method Fomicheva et al. (2020) relies on surface-form similarity, measuring the degree of word- or phrase-level overlap between generated texts. More advanced techniques propose various methods for taking into account the semantic similarity of the generated answers by hard or soft clustering (Lin et al., 2024).
Finally, given samples from the model, standard consistency-based methods compute a similarity matrix , where . Then, various statistics of are computed in order to estimate uncertainty (Lin et al., 2024), see details in Appendix D.2. Alternatively, one can focus on a particular output and compute uncertainty measures based on its similarities with other samples ; see Lin et al. (2024) for more details.
Interestingly, it was recently shown Wang and Holmes (2025); Daheim et al. (2025) that the latter approach has direct relation to MBR decoding. Building on the MBR framework, one can define the utility function as and derive the following consistency-based uncertainty score:
Then the Monte Carlo approximation of the minimum Bayes risk is given by:
where . In our ablation study, we show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix C.2).
A key strength of consistency-based techniques is that, by generating multiple samples and analyzing their semantic similarity, they can estimate empirical probabilities over meanings rather than individual answers. The main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of probabilities of specific answers.
@@ 3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures. There exist approaches that bridge the gap between information-based and consistency-based methods that show great promise but lack the fundamental base; see discussion in Section 4. In what follows, we present a family of Confidence and Consistency-based Approaches (CoCoA) for UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs via the unified MBR-based framework.
If we consider only the consistency measure, like the one in (8), we omit the information that is contained in the model confidence. Thus, we want the resulting uncertainty measure to explicitly consider both semantic consistency and model-based uncertainty. Let us consider the risk of the form:
@@
Then, the corresponding Bayes risk is and for an output sequence , the resulting uncertainty measure becomes:
@@
The resulting uncertainty measure integrates both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences that are both intrinsically uncertain (high ) and semantically inconsistent with the other samples (high ), while keeping it low for the opposite scenario; see Figure 2.
CoCoA Light. Computing the consistency-based uncertainty measure within requires multiple samples, leading to significant computational overhead. To address this, we propose to approximate the behavior of with a learned function. Our method closely matches the original sampling-based measure in quality while requiring only greedy output generation during inference, eliminating the need for additional sampling. Notably, this approximation can be learned without access to ground-truth labels, relying solely on the input data and the associated uncertainty values.
Thus, the CoCoA uncertainty measure that incorporates an approximation of the consistency uncertainty can be defined as:
@@
We note that a similar approach (Kossen et al., 2024) has been previously proposed to approximate semantic entropy using hidden states from the model.
@@ 4 Related Work
Information-based methods. These methods are one of the most commonly used. They quantify uncertainty by analyzing the probability distributions of tokens within a given output. The simplest of these methods, Maximum Sequence Probability (MSP), measures the probability of the most likely output sequence given a specific input. Perplexity is another common uncertainty measure: it gauges how well the model predicts each token in a sequence and is formally defined as the exponential of the mean negative log-likelihood of those tokens (Fomicheva et al., 2020). The Mean Token Entropy method evaluates token-level predictions across the entire sequence by computing the average entropy of the token probability distributions at each position (Fomicheva et al., 2020). While these methods provide useful uncertainty estimates, they do not address the broader uncertainty inherent in generative tasks, as they rely on a single sequence and fail to capture the diversity of possible outputs that the model could generate for the same input.
@@
Consistency-Based Methods. These methods estimate uncertainty by generating multiple sequences from the model’s output distribution and analyzing the variability among the sampled outputs. These methods are particularly relevant to NLP tasks, as they account for scenarios in which multiple plausible outputs may exist for a given input. Among the simplest consistency-based methods are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian Lin et al. (2024), which quantify uncertainty by how many distinct “meanings” the model produces. While effective at capturing global variation, they do not yield uncertainty scores for individual responses. To overcome this, the diagonal of the Degree Matrix was proposed to assess how similar each output is to the rest, enabling per-response uncertainty estimation Lin et al. (2024). A related approach, Lexical Similarity, operates by calculating the average similarity of words or phrases across every pair of responses in the sample Fomicheva et al. (2020).
@@
Information-Based Methods with Repeated Sampling. More recent methods attempt to reconcile output consistency with information-based signals—i.e., they not only look at how varied outputs are, but also how probable each output is under the model’s own generation process. For instance, Semantic Entropy (Kuhn et al., 2023) groups outputs into semantically homogeneous clusters—capturing distinct “meanings” that might just differ in phrasing—and calculates entropy across these clusters. SentenceSAR refines this idea by defining “relevance” as the sum of pairwise similarities between a given sentence and others weighted by each sentence’s model-assigned generation probability (Duan et al., 2024). SAR combines the sentence-level relevance of SentenceSAR with token-level probability adjustments, attempting a more fine-grained balance of semantic and token-level uncertainty (Duan et al., 2024). Semantic Density, Qiu and Miikkulainen (2024) evaluates uncertainty by evaluating how densely a model’s generated response is situated within the semantic space of all possible outputs, with lower density indicating higher uncertainty. One limitation of sampling-based approaches is their computational cost: generating multiple outputs from the model can be expensive. Recent work has sought to address this by learning to predict semantic uncertainty directly from hidden states, allowing for the approximation of semantic entropy without repeated sampling Kossen et al. (2024). Finally, BSDetector (Chen and Mueller, 2024) proposes an additive composition of self-reported confidence with observed consistency, however it relies on questionable ability of the model to assess it’s own confidence in text form. Another limitation is that it requires selection of a trade-off coefficient between confidence and consistency parts.
@@
While these probability-weighted measures can provide deeper insights, they also sometimes underperform in practice (Vashurin et al., 2025), especially when the model’s probability estimates are unreliable, or when important nuances of semantic diversity are lost in the weighting process. Proper balancing raw output consistency with generation likelihood remains an open problem.
@@
Minimum Bayes Risk for LLMs. Recent work has proposed incorporating uncertainty into Minimum Bayes Risk decoding by introducing a posterior over model parameters, enabling uncertainty-aware generation Daheim et al. (2025). Complementing this, a Bayesian framework for subjective uncertainty quantification was developed yielding improved calibration in generation tasks (Wang and Holmes, 2025).

### v5

# CoCoA: A Minimum Bayes Risk Framework Bridging Confidence and Consistency for Uncertainty Quantification in LLMs

@@ Abstract
Uncertainty quantification for Large Language Models (LLMs) encompasses a diverse range of approaches, with two major families being particularly prominent: (i) information-based, which estimate model confidence from token-level probabilities, and (ii) consistency-based, which assess the semantic agreement among multiple outputs generated using repeated sampling. While several recent methods have sought to combine these two paradigms to improve uncertainty quantification performance, they often fail to consistently outperform simpler baselines. In this work, we revisit the foundations of uncertainty estimation through the lens of Minimum Bayes Risk decoding, establishing a direct link between uncertainty and the optimal decision-making process of LLMs. Building on these findings, we propose CoCoA, a unified framework that integrates model confidence with output consistency, yielding a family of efficient and robust uncertainty quantification methods. We evaluate CoCoA across diverse tasks, including question answering, abstractive text summarization, and machine translation, and demonstrate sizable improvements over state-of-the-art uncertainty quantification approaches.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advances in information retrieval (zhu2023large), question answering (kwiatkowski-etal-2019-natural), machine translation (kocmi-federmann-2023-large), and a broad range of other NLP applications. As these models become an integral part of our everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which focuses on estimating the confidence of model predictions.
@@
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (malinin2020uncertainty; fomicheva-etal-2020-unsupervised). A key limitation of these methods is that they cannot account for semantic variability across multiple possible responses for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationship between the sampled outputs (fomicheva-etal-2020-unsupervised; lin2023generating), capturing uncertainty as objective variability of meaning among the sampled outputs.
Information-theoretic and consistency-based methods have complementary strengths. For this reason, recent state-of-the-art methods aimed to unify these approaches (kuhn2023semantic; duan-etal-2024-shifting). We follow this direction and propose a way of quantifying risk as a combination of these basic measures. This results in a family of efficient and robust UQ techniques. Our approach, illustrated in Figure 2, combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Recently, it has been argued (wangsubjective; daheimuncertainty) that UQ in NLP tasks can be viewed through the lens of the Minimum Bayes Risk (MBR) framework. Following this, we formulate our methods as particular forms of risk functions under the MBR approach.
@@
We propose a new way of quantifying risk that combines information-theoretic and consistency-based measures, and derive a family of Confidence and Consistency-based Approaches (CoCoA) to uncertainty quantification in LLMs111Our code is available publicly at https://github.com/stat-ml/llm_uncertainty_cocoa.
We show that the consistency component can be approximated by a learned function trained on unlabeled held-out set with a negligible loss of performance, eliminating the need for costly repeated sampling from the LLM. We call this variation of our method CoCoA Light.
We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and machine translation. Our experiments demonstrate sizable improvements in the reliability and the robustness of UQ compared to state-of-the-art methods.
@@ 2.1 Language Model Decoding
LLMs define a probabilistic output distribution for a given input sequence . The standard ways to obtain a particular output given include various variants of sampling and greedy decoding, where
An alternative approach is to use Minimum Bayes Risk (MBR) decoding (kumar2004minimum):
@@
with being a pairwise loss function.
The standard choice in the MBR literature is to take , where the so-called utility function represents some notion of similarity between generated sequences and . Below we discuss how various decoding strategies and corresponding risk functions lead to well-grounded definitions of uncertainty.
@@ 2.2 From Risk to Uncertainty
Bayesian uncertainty measures are well known to be strongly connected with minimum risks of various kinds (xu2022minimum; kotelevskii2025risk). Recently, this connection was revisited in the context of natural language generation (wangsubjective; daheimuncertainty), leading to the development of new MBR-based UQ methods. More precisely, one can consider the smallest achievable risk within or, equivalently, to maximum expected utility:
@@
Single Sequence Information-Based Methods. Information-based methods estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. Within the framework of MBR decoding, consider . In this case, the Bayes risk corresponds to the expected zero-one loss of decoding:
This leads to one of the simplest information-based uncertainty measures, Sequence Probability (SP):
Although widely used as an uncertainty measure across various applications, including natural language generation, its connection to statistical risk has only recently been explored in the UQ literature (kotelevskii2022nonparametric; kotelevskii2025risk; aichberger2024rethinking).
Several other measures fall into this category, including Perplexity and Mean Token Entropy (fomicheva-etal-2020-unsupervised); see Appendix D.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
@@
LLMs only give the probability of a specific answer, even though the same meaning can often be conveyed in multiple ways. To obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
@@
The reliability of such methods depends heavily on the underlying LLM being well-calibrated, which is a quality that is hard to define and harder to ensure.
@@
Semantic Consistency-Based Methods. The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs , where . Consistency-based UQ methods rely on the diversity of the answers sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require a probability distribution estimated by an LLM and can be applied in a black-box setting, where only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, consistency-based methods start from defining some similarity function between arbitrary LLM generations and . The value indicates the complete equivalence between and , and indicates that there is no similarity. The similarity could be computed in various ways. For instance, the Lexical Similarity method (fomicheva-etal-2020-unsupervised) relies on surface-form similarity, measuring the degree of word-level or phrase-level overlap between the generated texts. More advanced techniques propose various methods for taking into account the semantic similarity between the generated answers by means of hard or soft clustering (lin2023generating).
Finally, given samples from the model, standard consistency-based methods compute a similarity matrix , where . Then, various statistics of are computed in order to estimate the uncertainty (lin2023generating); see Appendix D.2 for more detail. Alternatively, one can focus on a particular output and compute uncertainty measures based on its similarity with other samples ; see (lin2023generating) for more detail.
Interestingly, it was recently shown (wangsubjective; daheimuncertainty) that the latter approach has direct relation to MBR decoding. Building on the MBR framework, one can define the utility function as and derive the following consistency-based uncertainty score:
Then the Monte Carlo approximation of the minimum Bayes risk is given by
where .
In our ablation study below, we show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all samples (see Appendix C.2).
A key strength of consistency-based techniques is that, by generating multiple samples and analyzing their semantic similarity, they can estimate empirical probabilities over meanings rather than over individual answers. Their main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of the probabilities of the specific answers.
@@ 3 CoCoA: Bridging Confidence and Consistency for Better Uncertainty Quantification
Both information-based and (semantic) consistency-based methods provide grounded and useful uncertainty quantification measures. There exist approaches that bridge the gap between information-based and consistency-based methods that show great promise but lack the fundamental base; see discussion in Section 4. In what follows, we present a family of Confidence and Consistency-based Approaches (CoCoA) for UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs via the unifying MBR-based framework.
If we only consider the consistency measure, e.g., as given in (8), we miss the information that is contained in the model confidence. Thus, we want the resulting uncertainty measure to explicitly consider both the semantic consistency and the model-based uncertainty. Let us consider a risk of the following form:
@@
Then, the corresponding Bayes risk is and for an output sequence , the resulting uncertainty measure becomes
@@
The resulting uncertainty measure integrates both global (semantic) and local (model-specific) uncertainty signals. It ensures that uncertainty is amplified for sequences that are both intrinsically uncertain (high ) and semantically inconsistent with respect to the other samples (high ), while keeping it low for the opposite scenario; see Figure 2 for an example.
CoCoA Light. Computing the consistency-based uncertainty measure within requires multiple samples, whoch poses significant computational overhead. To address this, we propose to approximate the behavior of with a learned function. Our method closely matches the original sampling-based measure in quality while requiring only greedy output generation during inference, eliminating the need for additional sampling. Notably, this approximation can be learned without access to ground-truth labels, relying solely on the input data and the associated uncertainty values.
Thus, the CoCoA uncertainty measure that incorporates an approximation of the consistency uncertainty can be defined as follows:
@@
We note that a similar approach (kossen2024semantic) has been previously proposed to approximate semantic entropy using the hidden states from the model.
@@ 4 Related Work
In this section, we review existing approaches to uncertainty quantification, as well as prior work on Minimum Bayes Risk in application to natural language generation.
Information-based methods.
@@
These methods are one of the most commonly used. They quantify uncertainty by analyzing the probability distributions of the tokens within a given output. The simplest of these methods, Sequence Probability (SP), measures the probability of the sequence given a specific input. Perplexity is another common uncertainty measure: it gauges how well the model predicts each token in a sequence and is formally defined as the exponential of the mean negative log-likelihood of those tokens (fomicheva-etal-2020-unsupervised). The Mean Token Entropy method evaluates the token-level predictions across the entire sequence by computing the average entropy of the token probability distributions at each position (fomicheva-etal-2020-unsupervised). While these methods provide useful uncertainty estimates, they do not address the broader uncertainty inherent in the generative tasks, as they rely on a single sequence and fail to capture the diversity of the possible outputs that the model could generate for the same input.
Consistency-Based Methods.
@@
These methods estimate uncertainty by generating multiple sequences from the model’s output distribution and analyzing the variability among the sampled outputs. These methods are particularly relevant to NLP tasks, as they account for scenarios in which multiple plausible outputs may exist for a given input. Among the simplest consistency-based methods are the Number of Semantic Sets and the Sum of Eigenvalues of the Graph Laplacian (lin2023generating), which quantify uncertainty by how many distinct “meanings” the model produces. While effective at capturing global variation, they do not yield uncertainty scores for individual responses. To overcome this, the diagonal of the Degree Matrix was proposed to assess how similar each output is to the rest, enabling per-response uncertainty quantification (lin2023generating). A related approach, Lexical Similarity, operates by calculating the average similarity of words or phrases across every pair of responses in the sample (fomicheva-etal-2020-unsupervised).
Information-Based Methods with Repeated Sampling.
@@
More recent methods attempt to reconcile output consistency with information-based signals—i.e., they not only look at how varied outputs are, but also how probable each output is under the model’s own generation process. For instance, Semantic Entropy (kuhn2023semantic) groups outputs into semantically homogeneous clusters—capturing distinct “meanings” that might just differ in phrasing—and calculates entropy across these clusters. SentenceSAR refines this idea by defining “relevance” as the sum of pairwise similarities between a given sentence and others weighted by each sentence’s model-assigned generation probability (duan-etal-2024-shifting). SAR combines the sentence-level relevance of SentenceSAR with token-level probability adjustments, attempting a more fine-grained balance of semantic and token-level uncertainty (duan-etal-2024-shifting). Semantic Density (qiu2024semantic) evaluates uncertainty by evaluating how densely a model’s generated response is situated within the semantic space of all possible outputs, with lower density indicating higher uncertainty. One limitation of sampling-based approaches is their computational cost: generating multiple outputs from the model can be expensive. Recent work has sought to address this by learning to predict semantic uncertainty directly from hidden states, allowing for the approximation of semantic entropy without repeated sampling (kossen2024semantic). Finally, BSDetector (chen-mueller-2024-quantifying) proposes an additive composition of self-reported confidence with observed consistency, however, it relies on ability of the model to assess its own confidence in text form. Another limitation is that it requires selection of a trade-off coefficient between confidence and consistency parts.
@@
While these probability-weighted measures can provide deeper insights, they also sometimes underperform in practice (vashurin2024benchmarkinguncertaintyquantificationmethods), especially when the model’s probability estimates are unreliable, or when important nuances of semantic diversity are lost in the weighting process. Proper balancing raw output consistency with generation likelihood remains an open problem.
Minimum Bayes Risk for LLMs.
@@
Minimum Bayes Risk (MBR) decoding, originally used in machine translation (Kumar2004MinimumBD), has recently been applied to LLMs by incorporating a posterior over model parameters to enable uncertainty-aware generation (daheimuncertainty). Complementing this, a Bayesian framework for subjective uncertainty quantification was developed yielding improved calibration in generation tasks (wangsubjective).